# myboxenhanced,colback=blue!5!white,colframe=blue!75!black,boxrule=0.6mm,fontupper=

### v1

@@ Abstract
Reinforcement Learning (RL), bolstered by the expressive capabilities of Deep Neural Networks (DNNs) for function approximation, has demonstrated considerable success in numerous applications. However, its practicality in addressing a wide range of real-world scenarios, characterized by diverse and unpredictable dynamics, noisy signals, and large state and action spaces, remains limited. This limitation stems from issues such as poor data efficiency, limited generalization capabilities, a lack of safety guarantees, and the absence of interpretability, among other factors. To overcome these challenges and improve performance across these crucial metrics, one promising avenue is to incorporate additional structural information about the problem into the RL learning process. Various sub-fields of RL have proposed methods for incorporating such inductive biases. We amalgamate these diverse methodologies under a unified framework, shedding light on the role of structure in the learning problem, and classify these methods into distinct patterns of incorporating structure. By leveraging this comprehensive framework, we provide valuable insights into the challenges associated with structured RL and lay the groundwork for a design pattern perspective on RL research. This novel perspective paves the way for future advancements and aids in the development of more effective and efficient RL algorithms that can potentially handle real-world scenarios better.
@@ 1 Introduction
Reinforcement Learning (RL) has contributed to a range of sequential decision-making and control problems like games (?), robotic manipulation (?), optimizing chemical reactions (?), and RNA folding (?). Most of the traditional research in RL focuses on designing agents that learn to solve a sequential decision problem induced by the inherent dynamics of a task, e.g., the differential equations governing the cart pole task in the classic control suite (?). However, their performance significantly degrades when even small aspects of the environment change (?, ?). Moreover, deploying RL agents for real-world learning-based optimization comes with additional challenges, such as complicated dynamics, intractable and computationally expensive state and action spaces, and noisy reward signals.
@@
Thus, research in RL has started to address these issues through methods that can generally be categorized into two dogmas (?): (i) Generalization:Methods developed to solve a broader class of problems where the agent is trained on a variety of different tasks and environments (?, ?). (ii) Deployability:Methods that are specifically engineered towards concrete real-world problems by incorporating additional aspects such as feature engineering, computational budget optimization, safety, etc. The intersection of generalization and deployability is particularly interesting since it covers a class of problems where we require methods to handle sufficient diversity in the task while being deployable for specific applications. To foster research in this area, ? (?) argue for a design-pattern oriented approach, where methods can be abstracted into patterns that are specialized to specific kinds of problems.
@@
However, the path to RL design patterns is hindered by gaps in our understanding of the relationship between the design decisions for RL methods and the properties of the problems they might be suited for. While decisions like using state abstractions for high-dimensional spaces seem obvious, decisions like using relational neural architectures for problems are not so obvious to a designer. One way to add principle to this process is to understand how to incorporate additional domain knowledge into the learning pipeline. A strong source of such domain knowledge is the structure present in the learning problem itself, including priors about the state space, the action space, the reward function, or the dynamics of the environment.
@@
Figure 1: Overview of our framework. Domain knowledge can generally be incorporated into an RL pipeline as side information and can be used to achieve improved performance across metrics such as Sample Efficiency, Generalization, Interpretability, and Safety. We discuss this process in Section 3. A particular source of side information is decomposability in a learning problem, which can be categorized into four archetypes along a spectrum - Latent, Factored, Relational, and Modular - explained further in Section 4.1. Incorporating side information about decomposability amounts to adding structure to a learning pipeline, and this process can be categorized into seven different patterns - Abstraction, Augmentation, Auxiliary Optimization, Auxiliary Model, Warehouse, Environment Generation, and Explicitly Designed - discussed further in Section 5.
@@
Contributions and Structure of the Paper. In this work, we take the first steps in this direction. Figure 1 shows a general overview of three elements of understanding the role of incorporating structure into a learning problem that we cover in this work. In Section 2, we describe the background and notation needed to formalize the relevant aspects of the RL problems. In Section 3, we introduce the notion of side information and define different additional metrics that can be addressed by incorporating side information into an RL pipeline. We then formulate structure as a particular kind of side information about decomposability in a problem in Section 4, and categorize decompositions in the literature into four major archetypes. In Section 5, we formulate seven patterns of incorporating structure into the RL learning process and provide an overview of each pattern by connecting it to the relevant surveyed literature. The framework developed in this work opens new avenues for research while providing a common reference point for understanding what kind of design decisions work under which situations. We discuss these aspects further in Section 6 for more concrete takeaways for both researchers and practitioners.
@@ 2.1 Markov Decision Processes
Sequential decision-making problems are usually formalized using the notion of a Markov Decision Process (MDP) (?, ?), which can be written down as a 5-tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,P,\rho\rangle$. At any timestep, the environment exists in a state $s\in\mathcal{S}$, with $\rho$ being the initial state distribution. The agent takes an action $a\in\mathcal{A}$ which transitions the environment to a new state $s^{\prime}\in\mathcal{S}$. The stochastic transition function governs the dynamics of such transitions $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, which takes the state $s$ and action $a$ as input and outputs a probability distribution over the next states $\Delta(.)$ from which the next state $s^{\prime}$ can be sampled. For each transition, the agent receives a reward $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, with $R\in\mathcal{R}$. The sequence $(s,a,r,s^{\prime})$ is called an experience.
The agent acts according to a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$, in a space of policies $\Pi$, that produces a probability distribution over actions given a state. This distribution is a delta distribution for deterministic policies, which leads to the policy outputting a single action. Using the current policy, an agent can repeatedly generate experiences, and a sequence of such experiences is also called a trajectory ($\tau$). In episodic RL, the trajectory consists of experiences collected over multiple episodes with environment resets, while in continual settings, the trajectory encompasses experiences collected over some horizon in a single episode. The rewards in $\tau$ can be accumulated into an expected sum called the return $G$, which can be calculated for any starting state $s$ as
$G(\pi,s)=\mathbb{E}_{(s_{0}=s,a_{1},r_{1},\dots)\sim\mathcal{M}}\bigg{[}\sum_{t=0}^{\infty}r_{t}\bigg{]}.$ (1)
For the sum in Equation 1 to be tractable, we either assume the horizon of the problem to be of a fixed length $T$ (finite-horizon return), i.e., the trajectory to terminate after $T$-steps, or we discount the future rewards by a discount factor $\gamma$ (infinite horizon return). Discounting, however, can also be applied to finite horizons. Solving an MDP amounts to determining the policy $\pi^{*}\in\Pi$ that maximizes the expectation over the returns of its trajectory. This expectation can be captured by the notion of the (state-action) value function $Q\in\mathcal{Q}$. Given a policy $\pi$, the expectation can be written recursively:
$Q^{\pi}(s,a)=\mathbb{E}_{s\sim\rho}\big{[}G_{t}\mid s,a\big{]}=\mathbb{E}_{s^{\prime}\sim\tau}\big{[}R(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[Q^{\pi}(s^{\prime},a^{\prime})]\big{]}.$ (2)
@@
$\pi^{*}\in\operatorname*{arg\,max}_{\pi\in\Pi}Q^{\pi}(s,a).$ (3)
We also consider partially observable MDPs (POMDPs), which model situations where the state is not fully observable. A POMDP is defined as a 7-tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},R,P,\xi,\rho\rangle$, where $\mathcal{S},\mathcal{A},R,P,\rho$ remain the same as defined above. Instead of observing the state $s\in\mathcal{S}$, the agent now has access to observation $o\in\mathcal{O}$ that is generated from the true state through an emission function $\xi:\mathcal{S}\to\mathcal{O}$. Thus, the observation takes the role of the state in the experience generation process, and the rest of the learning process can now be conditioned on $o$ instead of $s$.
@@ 2.2 Reinforcement Learning
The task of an RL algorithm is to interact with the MDP by simulating its transition dynamics $P(s^{\prime}\mid s,a)$ and reward function $R(s,a)$ and learn the optimal policy mentioned in Equation 3. In Deep RL, the policy is a Deep Neural Network (?) that is used to generate $\tau$. We can optimize such a policy by minimizing an appropriate objective $J\in\mathcal{J}$.
@@
RL methods can additionally be categorized based on the type of objective $J$. Methods that use a value function, and correspondingly the Temporal Difference error (?), to learn a policy fall into the category of Value-based RL. A key idea in these methods is bootstrapping, where they use a learned value estimate to improve the estimate of a state that precedes it. On-policy methods directly update the policy that generated the experiences, while Off-policy methods use a separate policy to generate experiences. Policy-Based Methods parameterize the policy directly and use the Policy Gradient (?, ?) to create $J$.
A central theme of research in practical RL methods focuses on approximating a global solution by iteratively learning one or more of the aforementioned quantities using supervised learning and function approximations. We use the notion of a pipeline to talk about different RL methods. Figure 2 shows the anatomy of an RL pipeline. The pipeline can be defined as a mathematical tuple $\Omega=\langle\mathcal{S},\mathcal{A},R,P,Q,\pi,\hat{\mathcal{M}},J,\mathcal{E}\rangle$, where all definitions remain the same as before. To solve an MDP, the pipeline operates on given an environment $\mathcal{E}$ by taking the state $s\in\mathcal{S}$ as input and producing an action $a\in\mathcal{A}$ as an output. The environment operates with the dynamics $P$ and a reward function $R$. The pipeline might generate experiences by directly interacting with $\mathcal{E}$, i.e., learning from experiences or by simulating a learned model $\hat{\mathcal{M}}$ of the environment. The optimization procedure encompasses the interplay between the current policy $\pi$, its value function $Q$, the reward $R$, and the learning objective $J$. With a slight abuse of notation, we refer to any of the components of a pipeline as $X$ and assume the space in which it exists as $\mathcal{X}$.

### v3

# Structure in Deep Reinforcement Learning: A Survey and Open Problems

@@ Abstract
Reinforcement Learning (RL), bolstered by the expressive capabilities of Deep Neural Networks (DNNs) for function approximation, has demonstrated considerable success in numerous applications. However, its practicality in addressing various real-world scenarios, characterized by diverse and unpredictable dynamics, noisy signals, and large state and action spaces, remains limited. This limitation stems from poor data efficiency, limited generalization capabilities, a lack of safety guarantees, and the absence of interpretability, among other factors. To overcome these challenges and improve performance across these crucial metrics, one promising avenue is to incorporate additional structural information about the problem into the RL learning process. Various sub-fields of RL have proposed methods for incorporating such inductive biases. We amalgamate these diverse methodologies under a unified framework, shedding light on the role of structure in the learning problem, and classify these methods into distinct patterns of incorporating structure. By leveraging this comprehensive framework, we provide valuable insights into the challenges of structured RL and lay the groundwork for a design pattern perspective on RL research. This novel perspective paves the way for future advancements and aids in developing more effective and efficient RL algorithms that can potentially handle real-world scenarios better.
@@ 1 Introduction
Reinforcement Learning (RL) has contributed to a range of sequential decision-making and control problems like games (?), robotic manipulation (?), and optimizing chemical reactions (?). Most of the traditional research in RL focuses on designing agents that learn to solve a sequential decision problem induced by the inherent dynamics of a task, e.g., the differential equations governing the cart pole task (?) in the classic control suite of OpenAI Gym (?). However, their performance significantly degrades when even minor aspects of the environment change (?, ?). Moreover, deploying RL agents for real-world learning-based optimization has additional challenges, such as complicated dynamics, intractable and computationally expensive state and action spaces, and noisy reward signals.
@@
Thus, research in RL has started to address these issues through methods that can generally be categorized into two dogmas (?): (i) Generalization:Methods developed to solve a broader class of problems where the agent is trained on various tasks and environments (?, ?). (ii) Deployability:Methods specifically engineered towards concrete real-world problems by incorporating additional aspects such as feature engineering, computational budget optimization, and safety. The intersection of generalization and deployability covers problems requiring methods to handle sufficient diversity in the task while being deployable for specific applications. To foster research in this area, ? (?) argue for a design-pattern oriented approach, where methods can be abstracted into patterns that are specialized to specific problems.
@@
However, the path to RL design patterns is hindered by gaps in our understanding of the relationship between the design decisions for RL methods and the properties of the problems they might be suited for. While decisions like using state abstractions for high-dimensional spaces seem obvious, decisions like using relational neural architectures for problems are not so apparent to a designer. One way to add principle to this process is to understand how to incorporate additional domain knowledge into the learning pipeline. The structure of the learning problem itself, including priors about the state space, the action space, the reward function, or the dynamics of the environment, is a vital source of such domain knowledge. While such methods have been research subjects throughout the history of RL (?), approaches that try to achieve this in Deep RL are scattered across the various sub-fields in the vast and disparate landscape of modern RL research. In this work, we take the first steps to amalgamate these approaches under our pattern-centric framework for incorporating structure in RL. Figure 1 shows a general overview of three elements of understanding the role of incorporating structure into a learning problem that we cover in this work.
@@
Figure 1: Overview of our framework. Side information can be used to achieve improved performance across metrics such as Sample Efficiency, Generalization, Interpretability, and Safety. We discuss this process in Section 4. A particular source of side information is decomposability in a learning problem, which can be categorized into four archetypes along a spectrum - Latent, Factored, Relational, and Modular - explained further in Section 5.1. Incorporating side information about decomposability amounts to adding structure to a learning pipeline, and this process can be categorized into seven different patterns - Abstraction, Augmentation, Auxiliary Optimization, Auxiliary Model, Warehouse, Environment Generation, and Explicitly Designed - discussed further in Section 6.
We refer to side information as additional knowledge not required to formulate the vanilla MDP. Incorporating structure entails utilizing side information about decomposability to improve sample efficiency, generalization, interpretability, and/or safety. To build an intuition about what we mean by this, consider the task of tailoring educational content to individual learners based on their preferences, learning pace, and mastery levels. An RL agent can select appropriate learning materials, activities, and assessments that best fit a learner’s current state and learning goals. Such a scenario is rich with structural properties and decompositions, such as learning styles or hidden skill proficiency of learners, distinct areas of knowledge within a learning program, relationships between knowledge areas and skills of learners, and modular content delivery mechanisms. While an MDP can potentially be formulated by treating the problem as a monolith, it does not need to be the most efficient solution. Instead, the problem can be formulated in various ways where prior knowledge about such decomposability can encode inductive biases into the RL agent. Prior knowledge about decompositions can additionally be discovered through auxiliary methods, such as Large Language Models (LLMs), that can analyze vast amounts of educational content, extracting key concepts, learning objectives, and difficulty levels. Incorporating side information into the learning pipeline, such as using the LLM to generate an intrinsic reward (?), can improve the speed of convergence of the RL agent, make it robust to variations in the problem and potentially help with making it safer and more interpretable.
Structure of the Paper.
@@
To better guide the reader, the paper is structured as follows: (i) In Section 2 we discuss the related works. We cover previous surveys on different areas in RL and previous works aimed at incorporating domain knowledge into RL methods. (ii) In Section 3, we describe the background and notation needed to formalize the relevant aspects of the RL problems. We additionally define the RL pipeline that we use in the later sections. (iii) In Section 4, we introduce side information and define the additional metrics that can be addressed by incorporating side information into an RL pipeline. (iv) In Section 5, we formulate structure as side information about decomposability and categorize decompositions in the literature into four archetypes on the spectrum of decomposability (?). Using these archetypes, we demonstrate how various problem formulations in RL fall into the proposed framework. (v) In Section 6, we formulate seven patterns of incorporating structure into the RL learning process and provide an overview of each pattern by connecting it to the relevant surveyed literature. We represent each pattern graphically as a plug-and-play modification to the RL pipeline introduced in Section 3. We additionally provide a literature survey for each pattern as a table and show possible research areas as empty spaces. (vi) In Section 7, we discuss how our framework opens new avenues for research while providing a common reference point for understanding what kind of design decisions work under which situations. We additionally summarize concrete takeaways for researchers and practitioners in various research areas in RL.
Scope of the Work.
Using structure has had a long history in RL, with early ideas already surfacing in the classical RL literature (?, ?, ?). Given the vast nature of this class of methods, we limit our focus in three ways: (i) we primarily cover Deep RL methods developed in the last ten years. While we discuss earlier works to establish the traditional underpinnings of our conceptual framework in Section 5, our survey in Section 6 only covers the later Deep RL literature; (ii) we only cover single-agent RL in this work. Multi-agent RL (MARL) (?) provides an additional dimension of decomposability, enabling additional patterns. While mathematically, certain aspects of such settings can be modeled equivalently by subsuming certain notions into the single-agent RL framework, we consider this field to have sufficient nuance and complexity to deserve a separate in-depth analysis and (iii) we do not cover theoretical work related to structure in RL. Such methods study how incorporating additional information and decompositions affect metrics such as learning complexity (?, ?), and could be potentially categorized into our suggested framework. However, we focus on empirical research because it is more widely studied.
2 Related Work
Multiple surveys have previously covered different areas in RL. However, none have covered the methods of explicitly and holistically incorporating structure in RL. In the following sections, we divide our literature research into surveys that tackle different problem settings, additional objectives, individual decompositions, and previous works incorporating domain knowledge into RL pipelines.
Different RL settings.
? (?) survey the field of Zero-Shot generalization and briefly discuss the need for more restrictive structured assumptions for their setting. While their survey argues for the requirement of similar assumptions, our work specifically lays out a framework that allows surveying approaches to utilize these assumptions. Additionally, our work is not limited to the setting of zero-shot generalization but covers additional areas of interpretability, safety, and sample efficiency in RL. ? (?) cover the field of Meta-RL and discuss the role of structure in Meta-Exploration, Transfer, and the POMDP formulation of Meta-Learning. However, their focus is on surveying the Meta-Learning setting and does not delve deeper into grounding what structure means, as is the case with our work. This is also the case with the survey of exploration methods in RL (?), where they argue for the need to choose the policy space to reflect prior information about the structure of the solution to ensure that the exploration behavior follows the same structure. Our framework grounds this idea in decomposition and argues for incorporating this information using one of many patterns.
Additional objectives.
Individual surveys have additionally covered the multiple objectives defined in Section 4. ? (?) provides a comprehensive review of the literature on safety in RL and divides the methods based on whether they modify the optimization criterion or the exploration process. We use their categorization to examine the correlation of patterns that use structural information for safety but cover additional objectives beyond safety. In a similar vein, ? (?) cover methods that add interpretability to the RL pipeline, which judges interpretability along the same axis as our work, namely, the definitions proposed by ? (?).
Grounding decompositions.
The assumptions of decomposability proposed in Section 5.1 utilize the spectrum of decomposability previously proposed by ? (?). We add to this framework by pinpointing four major decomposability archetypes on this spectrum, allowing us to build our categorization framework. Previous surveys have covered ideas related to these archetypes individually. ? (?) survey hierarchical methods that come under modular decompositions in our framework. ? (?) survey methods that leverage reasoning and declarative knowledge for sequential decision-making, including RL. A class of such methods falls under the relational decompositions in our framework.
Incorporating domain knowledge into RL.
Certain surveys have also been conducted on methods incorporating domain knowledge into RL. ? (?) survey methods that incorporate additional knowledge to tackle real-world deployment in RL. To this end, they categorize sources of knowledge into three types: (i) Scientific Knowledge, that covers empirical knowledge about the problem; (ii) World Knowledge, that covers an intuitive understanding of the problem that can be incorporated into the pipeline; and, (iii) Expert knowledge, available to experienced professionals in the form of experience. They formalize an RL pipeline and then look at methods incorporating this knowledge into different parts of the pipeline, such as problem representation, learning strategy, task structuring, and Sim2Real transfer. In addition to our scope focusing on domain knowledge about decomposability, our approach is source-agnostic. We focus on the specific part of the MDP on which structural assumptions are imposed and the nature of such assumptions. We categorize methods into patterns of incorporating these problems’ assumptions. Additionally, our patterns framework covers a broader range of methods that apply to more settings than Sim2Real.
The intersection of side information and patterns has previously been discussed by ? (?) and inspires our categorization as well. However, they predominantly discuss patterns for supervised and semi-supervised settings and mention trivial extensions to state representations in RL. Our formulation of patterns covers the RL pipeline more holistically by additionally looking at assumptions on components such as actions, transition dynamics, learned models, and previously learned skills. Moreover, our formulation solely focuses on different ways of biasing RL pipelines, holding little relevance for supervised and semi-supervised learning communities.
@@ 3.1 Markov Decision Processes
Sequential decision-making problems are usually formalized using the notion of a Markov Decision Process (MDP) (?, ?), which can be written down as a 5-tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,P,\rho\rangle$. At any timestep, the environment exists in a state $s\in\mathcal{S}$, with $\rho$ being the initial state distribution. The agent takes an action $a\in\mathcal{A}$ which transitions the environment to a new state $s^{\prime}\in\mathcal{S}$. The stochastic transition function governs the dynamics of such transitions $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, which takes the state $s$ and action $a$ as input and outputs a probability distribution over the following states $\Delta(.)$ from which the subsequent state $s^{\prime}$ can be sampled. For each transition, the agent receives a reward $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, with $R\in\mathcal{R}$. The sequence $(s,a,r,s^{\prime})$ is an experience.
The agent acts according to a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$, in a space of policies $\Pi$, that produces a probability distribution over actions given a state. This is a delta distribution for deterministic policies, which leads to the policy outputting a single action. Using the current policy, an agent can repeatedly generate experiences, and a sequence of such experiences is also called a trajectory ($\tau$):
$\tau=\{(s_{t},a_{t},r_{t},s_{t+1})\}_{t\in[t_{0},t_{T-1}]}\,\,\,\,\forall(s,a,r,s)\in\mathcal{S}\times\mathcal{A}\times R\times\mathcal{S}.$
In episodic RL, the trajectory consists of experiences collected over multiple episodes with environment resets. In contrast, in continual settings, the trajectory encompasses experiences collected over some horizon in a single episode. The rewards in $\tau$ can be accumulated into an expected sum called the return $G$, which can be calculated for any starting state $s$ as
$G(\pi,s)=\mathbb{E}_{(s_{0}=s,a_{1},r_{1},\dots)\sim\pi}\bigg{[}\sum_{t=0}^{\infty}r_{t}\bigg{]}.$ (1)
For the sum in Equation 1 to be tractable, we either assume the horizon of the problem to be of a fixed length $T$ (finite-horizon return), i.e., the trajectory to terminate after $T$-steps, or we discount the future rewards by a discount factor $\gamma$ (infinite horizon return). Discounting, however, can also be applied to finite horizons. Solving an MDP amounts to determining the policy $\pi^{*}\in\Pi$ that maximizes the expectation over the returns of its trajectory. This expectation can be captured by the (state-action) value function $Q\in\mathcal{Q}$. Given a policy $\pi$, the expectation can be written recursively:
$Q^{\pi}(s,a)=\mathbb{E}_{\pi}\big{[}\sum_{t=0}^{T}r_{t}\mid s_{0}=s,a_{0}=a\big{]}=\mathbb{E}_{\pi}\big{[}R(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[Q^{\pi}(s^{\prime},a^{\prime})]\big{]}.$ (2)
@@
$\pi^{*}\in\operatorname*{arg\,max}_{\pi\in\Pi}Q^{\pi}(s,a)\,\,\,\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (3)
We also consider Partially Observable MDPs (POMDPs), which model situations where the state is not fully observable. A POMDP is defined as a 7-tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},R,P,\xi,\rho\rangle$, where $\mathcal{S},\mathcal{A},R,P,\rho$ remain the same as defined above. Instead of observing the state $s\in\mathcal{S}$, the agent now has access to observation $o\in\mathcal{O}$ that is generated from the actual state through an emission function $\xi:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{O})$. Thus, the observation takes the state’s role in the experience generation process. However, solving POMDPs requires maintaining an additional belief since multiple $(s,a)$ can lead to the same $o$.
@@ 3.2 Reinforcement Learning
The task of an RL algorithm is to interact with the MDP by simulating its transition dynamics $P(s^{\prime}\mid s,a)$ and reward function $R(s,a)$ and learn the optimal policy mentioned in Equation 3. In Deep RL, the policy is a Deep Neural Network (?) that is used to generate $\tau$. We can optimize such a policy by minimizing an appropriate objective $J$.
@@
RL methods can additionally be categorized based on the type of objective $J$. Methods that use a value function, and correspondingly either Monte-Carlo estimates or Temporal Difference (TD) error (?), to learn a policy fall into the category of Value-based RL. A key idea in TD methods is bootstrapping, where they use a learned value estimate to improve the estimate of a state that precedes it. On-policy methods directly update the policy that generated the experiences, while Off-policy methods use a separate policy to generate experiences. Policy-Based Methods parameterize the policy directly and use the policy gradient theorem (?, ?) to create $J$.
A central research theme in practical RL methods focuses on approximating a global solution by iteratively learning one or more of the aforementioned quantities using supervised learning and function approximations. We use the notion of a pipeline to talk about different RL methods. Figure 2 shows the anatomy of an RL pipeline. The pipeline can be defined as a mathematical tuple $\Omega=\langle\mathcal{S},\mathcal{A},R,P,Q,\pi,\hat{\mathcal{M}},J,\mathcal{E}\rangle$, where all definitions remain the same as before. To solve an MDP, the pipeline operates on given an environment $\mathcal{E}$ by taking the state $s\in\mathcal{S}$ as input and producing an action $a\in\mathcal{A}$ as an output. The environment operates with the dynamics $P$ and a reward function $R$. The pipeline might generate experiences by directly interacting with $\mathcal{E}$, i.e., learning from experiences or by simulating a learned model $\hat{\mathcal{M}}$ of the environment. The optimization procedure encompasses the interplay between the current policy $\pi$, its value function $Q$, the reward $R$, and the learning objective $J$.